# On the Algorithmic Bias of Aligning Large Language Models with RLHF: Preference Collapse and Matching Regularization

### v1

@@ Abstract
Accurately aligning large language models (LLMs) with human preferences is crucial for informing fair, economically sound, and statistically efficient decision-making processes. However, we argue that reinforcement learning from human feedback (RLHF)—the predominant approach for aligning LLMs with human preferences through a reward model—suffers from an inherent algorithmic bias due to its Kullback–Leibler-based regularization in optimization. In extreme cases, this bias could lead to a phenomenon we term preference collapse, where minority preferences are virtually disregarded. To mitigate this algorithmic bias, we introduce preference matching (PM) RLHF, a novel approach that provably aligns LLMs with the preference distribution of the reward model under the Bradley–Terry–Luce/Plackett–Luce model. Central to our approach is a PM regularizer that takes the form of the negative logarithm of the LLM’s policy probability distribution over responses, which helps the LLM balance response diversification and reward maximization. Notably, we obtain this regularizer by solving an ordinary differential equation that is necessary for the PM property. For practical implementation, we introduce a conditional variant of PM RLHF that is tailored to natural language generation. Finally, we empirically validate the effectiveness of conditional PM RLHF through experiments on the OPT-1.3B and Llama-2-7B models, demonstrating a 29% to 41% improvement in alignment with human preferences, as measured by a certain metric, compared to standard RLHF.
@@ 1 Introduction
Large language models (LLMs) such as ChatGPT-4 (OpenAI, 2023) and Claude-3 Opus (Anthropic, 2024) have demonstrated impressive capabilities in tasks such as code generation, basic data analysis, elementary math problem solving, and reasoning (Bubeck et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023; Team et al., 2023). The widespread adoption of LLMs for tasks previously considered unlikely to be performed by machines reflects these advanced capabilities. For instance, a recent study revealed that a substantial proportion of machine learning conference reviews have been written or significantly modified by LLMs (Liang et al., 2024).
@@
As a result, LLMs are increasingly influencing decision-making processes across various domains (Eloundou et al., 2023; Duetting et al., 2023). It is, therefore, crucial to ensure that LLMs accurately reflect human preferences to promote fairness, mitigate the risk of biased outputs, and facilitate economically sound decisions (Arrow, 2012). Human preferences are inherently diverse due to factors such as cultural background, personal experiences, and contexts. It is natural for people from different backgrounds to exhibit varied preferences when asked to make comparisons. For example, Ouyang et al. (2022) and Bai et al. (2022) reported annotator-annotator and annotator-researcher disagreement rates of 37% and 23%, respectively, even for trained human labelers. Crucially, the relative proportions of these preferences play a critical role. For example, a restaurant chain deciding between adding a vegan dish or a spicy meat-based option to its menu might base its decision on whether customers have a 90% to 10% or a 20% to 80% preference for the two options. An LLM that fails to match human preferences could result in biased viewpoints on the tasks assigned to it. Consequently, the suggestions or decisions made by the LLM could suffer from fairness issues or, in this case, lead to undesirable economic outcomes.
@@
However, LLMs have been observed to exhibit significant biases when expressing preferences (Casper et al., 2023; Santurkar et al., 2023; Azar et al., 2024; Chakraborty et al., 2024). Casper et al. (2023) noted that LLMs tend to disproportionately favor more frequent items over less frequent ones. [^1] To illustrate, consider a scenario where humans prefer apples to oranges with a 70% to 30% ratio, while an LLM might express a 90% to 10% preference. This finding is in agreement with the observation by Santurkar et al. (2023) that opinion distributions generated by LLMs are often highly skewed towards the dominant viewpoints, often assigning over 99% probability to the dominant opinion and thus failing to capture the diversity of opinions within the population.
@@
In this paper, we aim to understand the origins of bias in LLMs from an algorithmic viewpoint and, more importantly, to develop a method for accurately aligning LLMs with human preferences with provable guarantees. To this end, it is instructive to examine reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Dong et al., 2024), the go-to solution for unleashing the power of LLMs by significantly simplifying the process of prompting for general tasks. [^2] To put this in perspective, LLMs had existed for at least five years—for example, GPT-3 was introduced in 2020—before the worldwide excitement surrounding the introduction of ChatGPT-3.5, which was enabled by applying standard RLHF to GPT-3. At its core, RLHF trains a reward model using the Bradley–Terry–Luce model (Bradley and Terry, 1952; Luce, 2012) to learn human preferences based on data from human labelers, and then fine-tunes the pretrained LLM using this reward model.
@@
First, we reveal an inherent bias in standard RLHF that arises from its Kullback–Leibler (KL) divergence-based regularization, which involves a pretrained LLM as a reference model for computing the divergence. We demonstrate that the reference model, arguably biased since it has not been aligned, can pass its bias to RLHF through the KL-based regularization. In certain cases, we theoretically show that this algorithmic bias can asymptotically amplify to extreme preference imbalances, such as 0% versus 100%, a phenomenon we term preference collapse. When preference collapse occurs, the LLM completely disregards minority opinions, which would lead to severe consequences in making fair and economical decisions. Furthermore, we show that merely replacing the KL divergence with a general $f$-divergence (Rényi, 1961) does not eliminate this algorithmic bias.
@@
To address the algorithmic bias and its extreme form, preference collapse, we introduce preference matching RLHF as the main contribution of this paper. This novel RLHF method provides a principled approach to precisely aligning LLMs with the reward model’s preferences, which comes with provable statistical guarantees. We start by considering a general function that regularizes the policy network, i.e., the probability distribution of responses given a prompt. By deriving a necessary condition for the regularizer to achieve preference matching in the form of an ordinary differential equation and solving it, we show that the regularization must equal the negative logarithm of the response probability distribution, up to certain less relevant terms. We further prove that this form of regularization is sufficient to match the reward model’s preferences. Taking the expectation, this regularization term becomes the Shannon entropy. Intuitively, in maximizing the reward plus the preference matching regularization, it seeks to strike a balance between reward maximization and diversifying the generation of responses. Moreover, we extend the differential equation-based strategy to obtaining preference matching regularization for a class of generalized Bradley–Terry–Luce/Plackett–Luce models.
@@
In comparison, a straightforward approach to addressing algorithmic biases is the early stopping of the fine-tuning process in standard RLHF, which appears to be one of OpenAI’s approaches based on publicly available information (Schulman, 2023). However, early stopping cannot entirely eliminate the bias in alignment since the “target” of standard RLHF is inherently biased, whereas our preference matching RLHF provably sets its global optimization solution to be unbiased.
@@
To examine the empirical performance of our preference matching RLHF, we introduce a variant of PM RLHF by conditioning PM RLHF on the set of responses that resemble natural language. We apply this conditional PM RLHF to fine-tune OPT-1.3B (Zhang et al., 2022) and Llama-2-7B (Touvron et al., 2023). In our experiments, we use proximal policy optimization (PPO) (Schulman et al., 2017) to train the model with our RLHF and the standard RLHF. To measure how well an RLHF approach does in terms of preference matching, we introduce a new measure called the preference matching divergence, defined as the KL divergence between the aligned model and the reward model. Using the standard RLHF approach, the preference matching divergence is 2.23 and 1.16 when fine-tuning Llama-2-7B and OPT-1.3B, respectively. Using our preference matching RLHF, the values of this divergence decrease to 1.57 and 0.68, marking improvements of approximately 29% and 41%, respectively.
@@
The remainder of the paper is structured as follows. Section 2 introduces the basics of LLMs and how standard RLHF operates on them. Section 3 introduces the concept of preference matching and obtains preference matching regularization by solving an ordinary differential equation. Section 4 extends preference matching RLHF to language domains, where we resolve a numerical challenge by introducing a conditional variant of our RLHF that is tailored to natural text generation. Section 5 formally argues why algorithmic bias emerges in standard RLHF and discusses its extreme case, preference collapse, explaining why it is inevitable in standard RLHF. [^3] We intentionally present Sections 3 and 4 before Section 5 to highlight the paper’s main results first. Experimental results are presented in Section 6. We conclude the paper with a discussion in Section 7.
@@ 1.1 Related Work
The observed bias in RLHF may originate from the assumption that complex human preferences can be captured by a single reward function. Chakraborty et al. (2024) challenged this assumption by providing a reward mismatch lower bound, arguing that diverse human preferences cannot be modeled by a single reward function. Similarly, Zhong et al. (2024) showed that a single reward function fails to balance the preferences of multiple heterogeneous individuals. To address this limitation, Wang et al. (2024) proposed directional preference alignment with multi-objective rewards to train LLMs for diverse user preferences.
@@
Training Algorithms.
Li et al. (2023b) demonstrated that PPO does not fully exploit the properties of RLHF in LLMs for preference alignment. Munos et al. (2023) introduced the Nash equilibrium of the preference model for aligning LLMs with human preferences, while Chen et al. (2024b) proposed a self-play-based approach to learn from supervised data as an alternative to supervised fine-tuning. Tang et al. (2024) discussed the performance gap between online and offline algorithms for alignment, and Ye et al. (2024) studied an online iterative RLHF algorithm with a general preference model.
A popular alternative to reward-model-based RLHF is direct preference optimization (DPO) (Rafailov et al., 2023), which directly fine-tunes LLMs on human preference data and has several notable variants (Liu et al., 2023; Azar et al., 2024; Chang et al., 2024; Gorbatovski et al., 2024; Rafailov et al., 2024; Yang et al., 2024). DPO is computationally more accessible since it bypasses the need for training a reward model. However, recent studies (Li et al., 2023a; Xu et al., 2024; Tajwar et al., 2024) suggest that DPO is inferior to reward-model-based RLHF in aligning LLMs. Li et al. (2023a); Xu et al. (2024) empirically argued that representation misspecification is the reason why DPO is inferior to RL-based methods such as PPO. Furthermore, the on-policy nature of fine-tuning using a reward model can help improve LLM performance, possibly due to the distribution shift between the offline preference dataset during training and online responses during evaluation (Tajwar et al., 2024).
@@ 2 Preliminaries
LLMs are large-scale deep learning architectures trained on extensive datasets (Vaswani et al., 2017; Kenton and Toutanova, 2019; Brown et al., 2020). They are built upon the Transformer architecture (Vaswani et al., 2017), which is designed to capture the context and relationships within textual sequences. This enables the models to extract and comprehend the meanings of words and phrases effectively.
[…]
@@ 2.1 Reinforcement Learning from Human Feedback
Let $x$ be a prompt given to an LLM and $y$ be the output response. In the second step, the SFT model is prompted with prompts $x$ to produce pairs of answers $(y_{1},y_{2})\sim\pi_{\textnormal{SFT}}$. For instance, consider the prompt $x$ being the question: “Please suggest a drink.” Two possible responses to this prompt are $y_{1}$ = “I would suggest coffee” and $y_{2}$ = “I would suggest tea”, respectively. After that, human labelers express their preferences for one of the two answers, labeled as $y_{\textnormal{w}}\succ y_{\textnormal{l}}$, where $y_{\textnormal{w}}$ and $y_{\textnormal{l}}$ denote the winner and loser among $(y_{1},y_{2})$, respectively.
@@
$\mathbb{P}(y_{1}\succ y_{2}|y_{1},y_{2},x)=\frac{\mathrm{e}^{r(x,y_{1})}}{\mathrm{e}^{r(x,y_{1})}+\mathrm{e}^{r(x,y_{2})}}.$ (2.1)
@@
where $\sigma(u):=1/(1+\mathrm{e}^{-u})$ is the logistic function. Hence, we omit the dependence of the reward function $r_{\theta}$ on $\theta$ because this paper does not discuss how the reward model is trained.
Extending the binary comparison settings of the BTL model to multi-comparison settings yields the Plackett–Luce (PL) model (Plackett, 1975; Luce, 2012), which also encompasses the BTL model. The PL model outputs the preference in the form of a ranking of different items, which will not be used in our paper. Therefore, we provide the standard assumption of the PL model in Assumption A.1 in the appendix. For simplicity, we consider an equivalent formulation of the PL model that assumes the preference of one response over others:
$\mathbb{P}(y|x)=\frac{\mathrm{e}^{r(x,y)}}{\sum_{y^{\prime}}\mathrm{e}^{r(x,y^{\prime})}},$
where the sum is over all possible responses to the prompt $x$.
@@
Let $\pi_{\phi}(y|x)$ be the probability distribution of the responses given a prompt $x$, where $\phi$ denotes the weights of the LLM. The goal of (unregularized) RLHF fine-tuning is to maximize the expected reward, i.e., $\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y)$. To mitigate over-optimization of the reward model, additional regularization should be added to the objective function, i.e.,
$\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y)-\text{Regularizer},$
where, unless specified otherwise, the expectation is taken over the randomness of $x$ and, conditional on $x$, the randomness of $y$ sampled from $\pi_{\phi}(\cdot|x)$. In practice, $x$ can be either sampled from a fixed database of prompts (Rafailov et al., 2023) or adaptively based on prior responses (Xiong et al., 2023). RLHF uses a KL penalty between the RLHF model and the reference model, which is the pretrained or SFT model, at each token as the regularizer. The loss function of regularized RLHF fine-tuning is
@@
where $\beta>0$ is a parameter controlling the deviation from the base reference policy $\pi_{\textnormal{ref}}$, namely the initial SFT model $\pi_{\text{SFT}}$. Above, the Kullback–Leibler divergence between $\pi_{\phi}(y|x)$ and $\pi_{\textnormal{ref}}(y|x)$ is defined as
$D_{\text{KL}}(\pi_{\phi}(y|x)\|\pi_{\textnormal{ref}}(y|x))=\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\big{[}\log(\pi_{\phi}(y|x))-\log(\pi_{\textnormal{ref}}(y|x))\big{]}.$
In practice, $\pi_{\phi}(y|x)$ is calculated token-wise, let $y=[T_{1},\cdots,T_{L}]$, it can be expressed in

### v2

@@ Abstract
Accurately aligning large language models (LLMs) with human preferences is crucial for informing fair, economically sound, and statistically efficient decision-making processes. However, we argue that the predominant approach for aligning LLMs with human preferences through a reward model—reinforcement learning from human feedback (RLHF)—suffers from an inherent algorithmic bias due to its Kullback–Leibler-based regularization in optimization. In extreme cases, this bias could lead to a phenomenon we term preference collapse, where minority preferences are virtually disregarded. To mitigate this algorithmic bias, we introduce preference matching (PM) RLHF, a novel approach that provably aligns LLMs with the preference distribution of the reward model under the Bradley–Terry–Luce/Plackett–Luce model. Central to our approach is a PM regularizer that takes the form of the negative logarithm of the LLM’s policy probability distribution over responses, which helps the LLM balance response diversification and reward maximization. Notably, we obtain this regularizer by solving an ordinary differential equation that is necessary for the PM property. For practical implementation, we introduce a conditional variant of PM RLHF that is tailored to natural language generation. Finally, we empirically validate the effectiveness of conditional PM RLHF through experiments on the OPT and Llama‑family models, demonstrating a 29% to 41% improvement in alignment with human preferences, as measured by a certain metric, compared to standard RLHF.
footnotemark: footnotetext: This paper has been accepted for publication in the Journal of the American Statistical Association.
@@ 1 Introduction
Large language models (LLMs) such as ChatGPT-4o (OpenAI, 2023) and Claude-3.7 Sonnet (Anthropic, 2024) have demonstrated impressive capabilities in tasks such as code generation, basic data analysis, elementary math problem solving, and reasoning (Bubeck et al., 2023; Chowdhery et al., 2023; Touvron et al., 2023; Bubeck et al., 2023; Team et al., 2023). The widespread adoption of LLMs for tasks previously considered unlikely to be performed by machines reflects these advanced capabilities. For instance, a recent study revealed that a substantial proportion of machine learning conference reviews had been written or significantly modified by LLMs (Liang et al., 2024).
@@
As a result, LLMs are increasingly influencing decision-making processes across various domains (Eloundou et al., 2023; Duetting et al., 2023). It is, therefore, crucial to ensure that LLMs accurately reflect human preferences to promote fairness, mitigate the risk of biased outputs, and facilitate economically sound decisions (Arrow, 2012). Human preferences are inherently diverse due to factors such as cultural background, personal experiences, and contexts. It is natural for people from different backgrounds to exhibit varied preferences when making comparisons. For example, Ouyang et al. (2022) and Bai et al. (2022) reported annotator–annotator and annotator–researcher disagreement rates of 37% and 23%, respectively, even for trained human labelers. Crucially, the relative proportions of these preferences play a critical role. For example, a restaurant chain deciding between adding a vegan dish or a spicy meat-based option to its menu might base its decision on whether customers have a 90% to 10% or a 20% to 80% preference for the two options. An LLM that fails to match human preferences could result in biased viewpoints on the tasks assigned to it. Consequently, the suggestions or decisions made by the LLM could suffer from fairness issues or, in this case, lead to undesirable economic outcomes.
@@
Indeed, LLMs have been observed to exhibit significant biases when expressing preferences (Casper et al., 2023; Santurkar et al., 2023; Azar et al., 2024; Chakraborty et al., 2024). Casper et al. (2023) noted that LLMs tend to disproportionately favor more frequent items over less frequent ones. This finding is in agreement with the observation by Santurkar et al. (2023) that opinion distributions generated by LLMs are often highly skewed towards the dominant viewpoints, often assigning over 99% probability to the dominant opinion and thus failing to capture the diversity of opinions within the population.
@@
In this paper, we aim to understand the origins of bias in LLMs from an algorithmic viewpoint and, more importantly, to develop a method for accurately aligning LLMs with human preferences with provable guarantees. To this end, it is instructive to examine reinforcement learning from human feedback (RLHF) (Ouyang et al., 2022; Dong et al., 2024), the go-to solution for unleashing the power of LLMs by significantly simplifying the process of prompting for general tasks. To put this in perspective, LLMs had existed for at least five years—for example, GPT-3 was introduced in 2020—before the worldwide excitement surrounding the introduction of ChatGPT-3.5, which was enabled by applying standard RLHF to GPT-3. At its core, RLHF trains a reward model using the Bradley–Terry–Luce model (Bradley and Terry, 1952; Luce, 2012) to learn human preferences based on data from human labelers, and then fine-tunes the pretrained LLM using this reward model.
@@
First, we reveal an inherent bias in standard Kullback–Leibler (KL) divergence-based RLHF, which involves a pretrained LLM as a reference model for computing the divergence. We demonstrate that the reference model, arguably biased since it has not been aligned, can pass its bias to RLHF through the KL-based regularization. In certain cases, we theoretically show that this algorithmic bias can asymptotically amplify to extreme preference imbalances, such as 0% versus 100%, a phenomenon we term preference collapse. When preference collapse occurs, the LLM completely disregards minority opinions, which would lead to severe consequences in making fair and economical decisions. Furthermore, we show that merely replacing the KL divergence with a general $f$-divergence (Rényi, 1961) does not eliminate this algorithmic bias.
@@
To address the algorithmic bias and its extreme form, preference collapse, we introduce preference matching RLHF as the main contribution of this paper. This novel RLHF method provides a principled approach to precisely aligning LLMs with the reward model’s preferences, which comes with provable statistical guarantees. We start by considering a general function that regularizes the policy network, i.e., the probability distribution of responses given a prompt. By deriving a necessary condition for the regularizer to achieve preference matching in the form of an ordinary differential equation and solving it, we show that the regularization must equal the negative logarithm of the response probability distribution, up to certain less relevant terms. We further prove that this form of regularization is sufficient to match the reward model’s preferences. Taking the expectation, this regularization term becomes the Shannon entropy. Intuitively, in maximizing the reward plus the preference matching regularization, it seeks to strike a balance between reward maximization and diversifying the generation of responses. Moreover, we extend the differential equation-based strategy to obtain preference matching regularization for a class of generalized Bradley–Terry–Luce/Plackett–Luce models.
@@
In comparison to our approach, a straightforward approach to addressing algorithmic biases is the early stopping of the fine-tuning process in standard RLHF, which appears to be one of OpenAI’s approaches based on publicly available information (Schulman, 2023). However, early stopping cannot entirely eliminate the bias in alignment since the “target” of standard RLHF is inherently biased, whereas our preference matching RLHF provably sets its global optimization solution to be unbiased.
@@
To examine the empirical performance of our preference matching RLHF, we introduce a variant of PM RLHF by conditioning PM RLHF on the set of responses that resemble natural language. We apply this conditional PM RLHF to fine-tune the OPT (Zhang et al., 2022) and Llama-family (Touvron et al., 2023) models. In our experiments, we use proximal policy optimization (PPO) (Schulman et al., 2017) to train the model with our RLHF and the standard RLHF. To measure how well an RLHF approach does in terms of preference matching, we introduce a new measure called the preference matching divergence, defined as the KL divergence between the aligned model and the reward model. Using the standard RLHF approach, the preference matching divergence is 2.23 and 1.16 when fine-tuning Llama-2-7B and OPT-1.3B, respectively. Using our preference matching RLHF, the values of this divergence decrease to 1.57 and 0.68, marking improvements of approximately 29% and 41%, respectively.
@@
The remainder of the paper is structured as follows. Section 2 introduces the basics of LLMs and how standard RLHF operates on them. Section 3 introduces the concept of preference matching and obtains preference matching regularization by solving an ordinary differential equation. Section 4 extends preference matching RLHF to language domains, where we resolve a numerical challenge by introducing a conditional variant of our RLHF that is tailored to natural text generation. Experimental results are presented in Section 5. We conclude the paper with a discussion in Section 6.
@@ 1.1 Related Work
The observed bias in RLHF may originate from the assumption that complex human preferences can be captured by a single reward function. Chakraborty et al. (2024) challenged this assumption by providing a reward mismatch lower bound, arguing that diverse human preferences cannot be modeled by a single reward function. Similarly, Zhong et al. (2024) showed that a single reward function fails to balance the preferences of multiple heterogeneous individuals. Xiao et al. (2025b) examined the inconsistency between reward learning and the principles of social choice theory. To address this limitation, Wang et al. (2024) proposed directional preference alignment with multi-objective rewards to train LLMs for diverse user preferences. Xu et al. (2024b) showed that methods such as RLHF can fail to yield desirable outcomes due to violations of the independence of irrelevant alternatives. Liu et al. (2025) and Shi et al. (2025) discussed the statistical impossibility of aligning with human preferences.
@@
Existing alignment approaches predominantly consider the average preference of human annotators, overlooking the rich diversity inherent in human preferences (Casper et al., 2023; Kaufmann et al., 2023; Li et al., 2025). Chakraborty et al. (2024) explored the factors contributing to this diversity, which often arises from various social and cultural backgrounds (Aroyo et al., 2023b). These factors include socio-demographic backgrounds, personal bias and context subjectivity, imperfect preferences, and linguistic ambiguity and missing context. Additional information on these factors is provided in the online Appendix.
@@ 2 Preliminaries
LLMs are large-scale deep learning architectures trained on extensive datasets (Vaswani et al., 2017; Kenton and Toutanova, 2019). They are built upon the Transformer architecture (Vaswani et al., 2017), which is designed to capture the context and relationships within textual sequences. This enables the models to extract and comprehend the meanings of words and phrases effectively.
The generation of sentences by LLMs relies on predicting the next token. Tokenization (Grefenstette, 1999) is the process of converting text into smaller units called tokens. Tokens can be words, subwords, characters, or other meaningful units. This process is a crucial step in natural language processing (NLP) and language modeling because it transforms raw text data into a format that can be processed by machine learning algorithms and models. For instance, the sentence “I just got a funky phone case!” can be tokenized into [“I”, “_just”, “_got”, “_a”, “_fun”, “ky”, “_phone”, “_case”, “!”], where “_” represents a space. Formally, tokenization can be expressed as $y=[T_{1},T_{2},T_{3},\ldots,T_{L}]$, where $L$ is the length of the sentence. Utilizing the sequence of preceding tokens, an LLM predicts the subsequent token. For example, the LLM assigns a probability to $p(T_{3}\mid\text{``I'', ``\_just''})$ across a range of possible next tokens such as [“_got”, “_took”, $\cdots$] from its vocabulary.
@@ 2.1 Reinforcement Learning from Human Feedback
Let $x$ be a prompt given to an LLM and $y$ be the output response. In the second step, the SFT model is prompted with prompts $x$ to produce pairs of answers $(y_{1},y_{2})\sim\pi_{\textnormal{SFT}}$. For instance, consider the prompt $x$ being the question: “Please suggest a drink.” Two possible responses to this prompt are $y_{1}$ = “I would suggest coffee” and $y_{2}$ = “I would suggest tea,” respectively. After that, human labelers express their preferences for one of the two answers, labeled as $y_{\textnormal{w}}\succ y_{\textnormal{l}}$, where $y_{\textnormal{w}}$ and $y_{\textnormal{l}}$ denote the winner and loser among $(y_{1},y_{2})$, respectively.
@@
$\mathbb{P}(y_{1}\succ y_{2}|y_{1},y_{2},x)=\frac{\exp(r(x,y_{1}))}{\exp(r(x,y_{1}))+\exp(r(x,y_{2}))}.$ (2.1)
@@
where $\sigma(u):=1/(1+\exp(-u))$ is the logistic function. Hence, we omit the dependence of the reward function $r_{\theta}$ on $\theta$ because this paper does not discuss how the reward model is trained.
@@
Let $\pi_{\phi}(y|x)$ be the probability distribution of the responses given a prompt $x$, where $\phi$ denotes the weights of the LLM. The goal of (unregularized) RLHF fine-tuning is to maximize the expected reward, i.e., $\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y)$. To mitigate over-optimization of the reward model, additional regularization should be added to the objective function, i.e., $\max_{\phi}\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}r(x,y)-\text{Regularizer},$ where, unless specified otherwise, the expectation is taken over the randomness of $x$ and, conditional on $x$, the randomness of $y$ sampled from $\pi_{\phi}(\cdot|x)$. In practice, $x$ can be either sampled from a fixed database of prompts (Rafailov et al., 2023) or adaptively based on prior responses (Xiong et al., 2023). RLHF uses a KL penalty between the RLHF model and the reference model, which is the pretrained or SFT model, at each token as the regularizer. The loss function of regularized RLHF fine-tuning is
@@
where $\beta>0$ is a parameter controlling the deviation from the base reference policy $\pi_{\textnormal{ref}}$, namely the initial SFT model $\pi_{\text{SFT}}$. Above, the Kullback–Leibler divergence between $\pi_{\phi}(y|x)$ and $\pi_{\textnormal{ref}}(y|x)$ is defined to be $\mathbb{E}_{y\sim\pi_{\phi}(\cdot|x)}\big{[}\log(\pi_{\phi}(y|x))-\log(\pi_{\textnormal{ref}}(y|x))\big{]}$. In practice, $\pi_{\phi}(y|x)$ is calculated token-wise, let $y=[T_{1},\cdots,T_{L}]$, it can be expressed in